Reports 1-1 of 1 Clear search Modify search
OBS (General)
Tomaru and Ohino - 1:04 Sunday 12 April 2020 (14117) Print this report
Shift Report at 5pm-25pm, 11 Apr., 2020
Interferometer lock was almost stable during our observation term. From around 7pm, the ITMY pitch started a bit oscillating and interferometer could not keep stable lock. So we made interferometer down once, Details logs are following;
7pm: The ITMY pitch started oscillation.
7:34pm: The Interferometer had locked-loss
7:42pm: Observation was restarted automatically.
7:51pm: Lock-loss again. Since ITMY pitch oscillation is a bit large, once we sent "down" command and waited to reduce it.
8:17pm: Observation restart
8:55pm: lock-loss again. We decided to do alignment by "PRFPMI_LSC_locked_RF"
9:27pm: Observation restarted
0:45am: Lock-loss due to earthquake. Once we did "Down" to wait to decrease excitation of VIS.
1:00: Requested "PRFPMI_LSC_locked_RF"
Search Help